OpenStudy (anonymous):

@mathmath333 Can you help me? I need to find how many fl oz is in 6 7/8 c. (Metric system)

OpenStudy (mathmath333):

is it \(6\cfrac{7}{8}\) or \(\cfrac{67}{8}\)

OpenStudy (mathmath333):

\(\Large \begin{align} \color{black}{1 c =8 ~~fl ~~~oz}\end{align}\)

OpenStudy (anonymous):

A) 110 fl oz B) 48 fl oz C) 55 fl oz D) 55/64 fl oz

OpenStudy (anonymous):

I don't know if my school is wrong or if we're just not getting it but no one has been able to help me and it really sucks.

OpenStudy (mathmath333):

oh \(\Large \begin{align} \color{black}{6\cfrac{7}{8}\\~\\ =\dfrac{6\times 8+7}{8}\\~\\ =\dfrac{55}{8}~~c\\~\\ =\dfrac{55}{8}\times 8 ~~fl~~oz} \end{align}\)

OpenStudy (anonymous):

Omg ty so so much!!!